OpenStudy (anonymous):

Find the missing coefficient in the equation of the line that passes through the given point.

OpenStudy (anonymous):

whats the given point?

OpenStudy (anonymous):

3x+By=-1 (2,7)

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

3*2+B*7=-1 6+B*7=-1 6+(-1*7) B=-1

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

medal please